# ALOE: Action-Level Off-Policy Evaluation for Vision-Language-Action Model Post-Training

### v1

@@ Abstract
We study how to improve large foundation vision-language-action (VLA) systems through online reinforcement learning (RL) in real-world settings. Central to this process is the value function, which provides learning signals to guide VLA learning from experience. In practice, the value function is estimated from trajectory fragments collected from different data sources, including historical policies and intermittent human interventions. Estimating the value function of current behavior quality from the mixture data is inherently an off-policy evaluation problem. However, prior work often adopts conservative on-policy estimation for stability, which avoids direct evaluation of the current high-capacity policy and limits learning effectiveness. In this paper, we propose ALOE, an action-level off-policy evaluation framework for VLA post-training. ALOE applies chunking-based temporal-difference bootstrapping to evaluate individual action sequences instead of predicting final task outcomes. This design improves effective credit assignment to critical action chunks under sparse rewards and supports stable policy improvement. We evaluate our method on three real-world manipulation tasks, including smartphone packing as a high-precision task, laundry folding as a long-horizon deformable-object task, and bimanual pick-and-place involving multi-object perception. Across all tasks, ALOE improves learning efficiency without compromising execution speed, showing that off-policy RL can be reintroduced in a reliable manner for real-world VLA post-training. Videos and additional materials are available at our project website.
@@ IV-A Off-Policy Critic Estimation under Real-World Data Collection
To mitigate these challenges, we adopt TD learning [42], which bootstraps from subsequent value estimates and allows value information to propagate backward through time. Starting from the Bellman equation,
$Q^{\pi}(s,a,\ell)=r+\gamma\,\mathbb{E}_{s^{\prime}\sim p,\,a^{\prime}\sim\pi}\left[Q^{\pi}(s^{\prime},a^{\prime},\ell)\right],$ (5)
@@
In regions of the state-action space with insufficient observations, we find that traditional TD-based methods may extrapolate aggressively and produce overestimated action values. This effect is amplified in safety-critical manipulation tasks, where exploration is constrained and data diversity is inherently limited. To mitigate the overestimation arising from uncertainty, we employ a pessimistic ensemble of $K$ action-value functions $\{Q_{\phi_{i}}\}_{i=1}^{K}$. Each critic is trained using the same chunked TD target in Eq. (7), by minimizing the objective
@@
where $y^{(h)}_{t}$ denotes the chunked TD target. At policy update time, we form a conservative estimate
@@ IV-B Advantage-Weighted Policy Improvement for VLA Policies
Given a transition, we evaluate the quality of the data action relative to the current policy $\pi$ by defining the advantage
@@
where the constraint $\epsilon$ is implicitly controlled by the temperature parameter $\beta$ (distinct from the clipping parameter $\epsilon_{\mathrm{clip}}$ in Eq. (11)).
@@ IV-C Implementation Details
$(s_{t},\mathbf{a}_{t:t+h},\ell)\sim\mathcal{D},\quad\eta\sim\mathcal{U}[0,1],\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (15)
with the noised action defined as $\tilde{\mathbf{a}}_{t:t+h}=\eta\mathbf{a}_{t:t+h}+(1-\eta)\boldsymbol{\epsilon}$. Intuitively, maximizing the weighted policy log likelihood corresponds to minimizing weighted flow-matching loss for the action chunks under the same inputs. The state input to the actor and critic is $s_{t}=[X^{1}_{t},X^{2}_{t},X^{3}_{t},q_{t}]$, where $X^{i}_{t}$ denotes the image from the $i$-th camera and $q_{t}$ denotes the robot proprioceptive observation. More training details and parameter settings can be found on Appendix -C
@@ V Experiments
In this section, we evaluate ALOE on a set of real-world, robotic manipulation tasks designed to stress long-horizon reasoning, precise action selection, and robustness under human-in-the-loop data collection. Specifically, we consider three representative tasks: Pack Smart Phone, Folding Laundry, and Product Sorting. These tasks require a combination of fine-grained manipulation, multi-stage task composition (e.g., pick–place–push), and reasoning over visually ambiguous intermediate states. All experiments are conducted on a real robotic platform, illustrated in Fig. 2. More experimental implementation can be found on Appendix -B

### v2

@@ Abstract
We study how to improve large foundation vision-language-action (VLA) systems through online reinforcement learning (RL) in real-world settings. Central to this process is the value function, which provides learning signals to guide VLA learning from experience. In practice, the value function is estimated from trajectory fragments collected from different data sources, including historical policies and intermittent human interventions. Estimating the value function of current behavior quality from the mixture data is inherently an off-policy evaluation problem. However, prior work often adopts conservative on-policy estimation for stability, which avoids direct evaluation of the current high-capacity policy and limits learning effectiveness. In this paper, we propose ALOE, an action-level off-policy evaluation framework for VLA post-training. ALOE applies chunking-based temporal-difference bootstrapping to evaluate individual action sequences instead of predicting final task outcomes. This design improves effective credit assignment to critical action chunks under sparse rewards and supports stable policy improvement. We evaluate our method on three real-world manipulation tasks, including smartphone packing as a high-precision task, laundry folding as a long-horizon deformable-object task, and bimanual pick-and-place involving multi-object perception. Across all tasks, ALOE improves learning efficiency without compromising execution speed, showing that off-policy RL can be reintroduced in a reliable manner for real-world VLA post-training. Videos and additional materials are available at our project website https://rooshy-yang.github.io/aloe.
@@ IV-A Off-Policy Critic Estimation under Real-World Data Collection
To mitigate these challenges, we adopt TD learning [42], which bootstraps from subsequent value estimates and allows value information to propagate backward through time. Starting from the Bellman equation:
$Q^{\pi}(s,a,\ell)=r+\gamma\,\mathbb{E}_{s^{\prime}\sim p,\,a^{\prime}\sim\pi}\left[Q^{\pi}(s^{\prime},a^{\prime},\ell)\right].$ (5)
@@
In regions of the state-action space with insufficient observations, we find that traditional TD-based methods may extrapolate aggressively and produce overestimated action values. This effect is amplified in safety-critical manipulation tasks, where exploration is constrained and data diversity is inherently limited. To mitigate the overestimation arising from uncertainty, we employ a pessimistic ensemble of $K$ action-value functions $\{Q_{\phi_{i}}\}_{i=1}^{K}$. Each critic is trained using the same chunked TD target in Eq. (7), by minimizing the objective:
@@
where $y^{(h)}_{t}$ denotes the chunked TD target. At policy update time, we form a conservative estimate:
@@ IV-B Advantage-Weighted Policy Improvement for VLA Policies
Given a transition, we evaluate the quality of the data action relative to the current policy $\pi$ by defining the advantage:
@@
The constraint $\epsilon$ is implicitly controlled by the temperature parameter $\beta$.
@@ IV-C Implementation Details
$(s_{t},\mathbf{a}_{t:t+h},\ell)\sim\mathcal{D},\quad\eta\sim\mathcal{U}[0,1],\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (15)
where the noised action defined as $\tilde{\mathbf{a}}_{t:t+h}=\eta\mathbf{a}_{t:t+h}+(1-\eta)\boldsymbol{\epsilon}$. Intuitively, maximizing the weighted policy log likelihood corresponds to minimizing weighted flow-matching loss for the action chunks under the same inputs. The state input to the actor and critic is $s_{t}=[X^{1}_{t},X^{2}_{t},X^{3}_{t},q_{t}]$, where $X^{i}_{t}$ denotes the image from the $i$-th camera and $q_{t}$ denotes the robot proprioceptive observation. More training details and parameter settings can be found on Appendix -C.
@@ V Experiments
In this section, we evaluate ALOE on a set of real-world, robotic manipulation tasks designed to stress long-horizon reasoning, precise action selection, and robustness under human-in-the-loop data collection. Specifically, we consider three representative tasks: Pack Smart Phone, Folding Laundry, and Product Sorting. These tasks require a combination of fine-grained manipulation, multi-stage task composition (e.g., pick–place–push), and reasoning over visually ambiguous intermediate states. All experiments are conducted on a real robotic platform, illustrated in Fig. 2. More experimental implementation can be found on Appendix -B.